The difference of two numbers is 16. The second is 9 less than 2 times the first. What are the two numbers?
Include the 2 equations and your work for solving using equations.

Answer by stanbon(75887)

The difference of two numbers is 16. The second is 9 less than 2 times the first. What are the two numbers?
----------------
x-y = 16
y = 2x-9
===========
Solve by substitution:
x-(2x-9) = 16
-x+9 = 16
x = -7
---------
Substitute to solve for y:
y = 2*-7-9
y = -23
==============
Cheers,
Stan H.